if y(t) = A sin (wt+Ø) What will be the equation for the velocity and acceleration?


1
Expert's answer
2020-11-30T14:53:24-0500
"v(t)=\frac{dy}{dt}=w\cos{(wt+\theta)}\\a(t)=\frac{dv}{dt}=-w^2\sin{(wt+\theta)}"